# Aligning Text to Image in Diffusion Models is Easier Than You Think

### v4

@@ Abstract
While recent advancements in generative modeling have significantly improved text-image alignment, some residual misalignment between text and image representations still remains. Some approaches address this issue by fine-tuning models in terms of preference optimization, etc., which require tailored datasets. Orthogonal to these methods, we revisit the challenge from the perspective of representation alignment—an approach that has gained popularity with the success of REPresentation Alignment (REPA) [45]. We first argue that conventional text-to-image (T2I) diffusion models, typically trained on paired image and text data (i.e., positive pairs) by minimizing score matching or flow matching losses, is suboptimal from the standpoint of representation alignment. Instead, a better alignment can be achieved through contrastive learning that leverages existing dataset as both positive and negative pairs. To enable efficient alignment with pretrained models, we propose SoftREPA—a lightweight contrastive fine-tuning strategy that leverages soft text tokens for representation alignment. This approach improves alignment with minimal computational overhead by adding fewer than 1M trainable parameters to the pretrained model. Our theoretical analysis demonstrates that our method explicitly increases the mutual information between text and image representations, leading to enhanced semantic consistency. Experimental results across text-to-image generation and text-guided image editing tasks validate the effectiveness of our approach in improving the semantic consistency of T2I generative models. Project Page: https://softrepa.github.io/
@@ 3.1 Contrastive T2I Alignment Loss
Here, $\tau(t)$ represents time scheduling parameter. One might consider using the denoising score matching loss, $-\mathbb{E}_{t,{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}]$, as a similarity measure instead of relying on its logit value. However, our experiments revealed that this formulation can introduce instability during training due to the unbounded nature of the similarity measure. To address this, SoftREPA introduce an exponential function to constrain the logit values, effectively stabilizing the training process.
@@ 3.2 The Soft Tokens
To distill the contrastive score matching objective above, we introduce a learnable soft token ${\bm{s}}$, which is the only trainable parameter while pretrained model remains frozen. In particular, as illustrated in fig. 2, we introduce a set of learnable soft tokens that vary between layers and are indexed by time $t$. The soft token at layer $k$ is defined using an embedding function with the number of soft tokens as $m$:
@@
Then, the learnable token ${\bm{s}}$ is optimized by minimizing $\mathcal{L}_{\text{SoftREPA}}({\bm{s}})$. During inference, the trained soft tokens are leveraged with fixed text representation from the text encoder across layers and timesteps. The detailed procedure during the image generation process based on MM-DiT architecture is described in algorithm 1 from appendix B.
@@ 3.3 Relationship with Mutual Information
where the second equality follows from the definition of conditional expectation. Song et al. [35] and Kong et al. [18] further show that assuming the diffusion model is an optimal denoiser, the conditional likelihood $p_{\theta}({\bm{x}}|{\bm{y}})$ can be described as (see appendix A):
[…]
@@
Human Preference Text Alignment Image Quality PM (GB) Latency (sec/image)
@@
Table 1: Quantitative evaluation of T2I generation with learnable soft tokens on SD1.5, SDXL, and SD3. Generation quality is evaluated on the COCO-val 5K [22] and GenEval [9] benchmark. Peak GPU memory(PM) usage and per-image latency are measured to assess computational efficiency. ImageReward, CLIP, HPS, and LPIPS are scaled by ×$\times 10^{2}$.
@@
Table 2: Quantitative evaluation of image editing performance with the use of soft tokens on PnP [37], MasaCtrl [4], Direct inversion [14], and FlowEdit [19]. FlowEdit is based on SD3, and others are based on SD1.5 architecture. The editing methods are evaluated on PIEBench [14], DIV2K [1], and Cat2Dog(C2D). ImageReward, CLIP, HPS, LPIPS, and SSIM are scaled by ×$\times 10^{2}$ and Distance is scaled by ×$\times 10^{3}$.
@@ 4 Experiments
Implementation details. In our experiments, to confirm the flexibility of SoftREPA, we utilize various open-source T2I diffusion models, including Stable Diffusion 1.5, Stable Diffusion XL, and Stable Diffusion 3 to evaluate text and image alignment in both image generation and text-based image editing tasks. For the training of softs token, the batch size was set to 16 with less than 30,000 iterations using two A100 GPUs. For Stable Diffusion 3, we set the length of the soft token to 4 and used 5 layers to attach the soft tokens for further experiments. In the case of Stable Diffusion 1.5 and Stable Diffusion XL, soft tokens are applied only to the Down or Middle block layers of the UNet. Refer to further implementation details in appendix B.
@@
We conducted text-to-image generation experiments on SD1.5, SDXL, and SD3, training soft tokens using the text-image paired COCO dataset [22]. The implementation details are provided in appendix B. To evaluate the generated images, we assessed human preference scores [44, 17], text-image alignment [47, 43], and image quality [46, 33] on COCO-val 5K dataset. Furthermore, we conducted evaluation on GenEval [9], compared with RankDPO [16], which is fine-tuned via preference optimization. As shown in table 1, in COCO dataset, the proposed method generally outperforms baseline approaches in both diffusion and rectified flow models regardless of the model architecture. Additionally, in GenEval, SD3 with SoftREPA mostly outperforms RankDPO [16]. The drop in the counting metric can be primarily attributed to the tendency of diffusion models to interpret text alignment in a way that encourages the generation of multiple instances of the referenced object. Further analysis and potential mitigation strategies are discussed in appendix D. In fig. 3, detailed textual descriptions are more accurately reflected in the generated images, as demonstrated in the qualitative results.
@@
In table 1, we report the peak GPU memory usage (PM) and per-image latency for models with and without soft tokens, including SD1.5, SDXL, and SD3. All models were evaluated in float16 precision. The use of soft tokens introduces no change in peak memory and results in only a negligible increase in latency.
@@
To further demonstrate the improved text-image alignment achieved by our method, we conducted text-guided image editing experiments using PnP [37], MasaCtrl [4] with DDIM [34], Direct Inversion [14], and FlowEdit [19]. FlowEdit employs SD3, while the other methods use SD1.5. For evaluation, we used PIEBench [14], DIV2K [1], and Cat2Dog datasets. For DIV2K editing dataset, we selected 800 high-quality images and generated detailed source and target captions using LLaVA [24]; dataset construction details are provided in appendix C. In table 2, the quantitative results of human preference, text alignment, and structure distance show consistent enhancements with the use of soft tokens without compromising structure and background fidelity.
Considering that FlowEdit performs text-based image editing, where performance heavily depends on text-image alignment, this result indicates improved efficiency in modeling the joint text-image distribution using the proposed method. Furthermore, our method with soft tokens achieves more efficient editing by reducing the number of editing steps to 30, compared to 33 in FlowEdit. As shown in fig. 6, incorporating soft tokens across various CFG scales consistently enhances text-image alignment and image quality. Figure 4 further demonstrates that performance has been improved on both single and multi-concept editing compared to vanilla SD3. Additional results and details are provided in appendix F.
@@
We conducted an ablation study on SD3 using the COCO val dataset [22] to evaluate the effects of the number of layers using soft text tokens and token length on generation quality (fig. 6). A value of 0 corresponds to vanilla SD3, and the number of layers is counted from the first. Comparisons were made using either a fixed token length of 8 or a fixed depth of 5 layers. Our findings indicate that incorporating soft tokens beyond layer $7$ severely degrades image generation quality. The optimal performance, aligned with human preferences (ImageReward [44]), was observed when soft tokens were applied within layers $2\sim 5$. This aligns with the architecture, where early layers handle text-image alignment and later ones enhance fidelity [26]. Notably, CLIP scores continued to improve as more layers used soft tokens, indicating stronger adherence to text prompts, while overall image quality remained stable. Supporting qualitative results are provided in appendix G. For token length, using more than 8 tokens negatively impacted quality, suggesting overfitting to the fine-tuning data. Shorter lengths (1–4 tokens) preserved perceptual similarity (LPIPS), while moderate lengths (4–8 tokens) significantly improved text-image alignment (CLIP and ImageReward). Qualitative examples are shown in appendix G.
@@ Appendix A Information Theoretical Analysis of Diffusion Models
[…]
@@
[…]
@@ Appendix B Implementation Details
The hyperparameters used during training are listed in table 3. For Stable Diffusion 3, the soft tokens are optimized solely using the contrastive score matching loss ($L_{\text{SoftREPA}}$). In contrast, for Stable Diffusion 1.5 and Stable Diffusion XL, optimization combines the contrastive score matching loss with a small weighting of the denoising score matching loss. We observed that initializing the soft tokens with a random distribution led to performance degradation in Stable Diffusion 1.5. To address this, we initialized the soft tokens using unconditional text embeddings. The code is publicly available at https://github.com/softrepa/SoftREPA.
@@
A detailed description of the image generation algorithm on MM-Dit is provided in algorithm 1. At each layer, a distinct set of soft tokens is prepended to the text features and used exclusively within that layer. These soft tokens are not carried over to subsequent layers; instead, new soft tokens are introduced or omitted as appropriate. Their primary role is to guide the text tokens toward better text-image alignment, particularly during the early stages of the model.
@@
For the memory efficiency metrics reported in table 1, values were measured by averaging results over 50 runs with an A100 GPU. Image resolution was set to $512\times 512$ for SD1.5 and $1024\times 1024$ for SDXL and SD3.
Algorithm 1 Image Generation with Soft Tokens in MM-DiT
@@
4:Number of layers $N$, Threshold layer $L$
@@
18:         $\bm{H}_{\text{text}}^{(k,t)}\leftarrow\hat{\bm{H}}_{\text{text}}^{(k,t)}[-n:,:]$triangleright Drop soft tokens
@@ Appendix D Discussion on counting metric
As shown in section 4, our model demonstrates notable improvements in generating single and multiple objects, as well as in color and position-aware synthesis. However, it exhibits a relative decline in accurately generating the specified number of objects described in the text prompts. We hypothesize that this drop in the counting metric stems from the soft tokens excessively emphasizing textual cues, which can lead to overgeneration of object instances. To address this, we limited the use of soft tokens to the early layers of the model. As presented in table 4, applying soft tokens only to layers $1\sim 2$ preserves strong overall performance while mitigating the degradation in counting accuracy observed when soft tokens are applied across layers $1\sim 5$.
@@
To further enhance counting fidelity, we trained SoftREPA with a lightweight object-counting loss. Specifically, we employ a mean squared error (MSE) loss between the predicted object count, which is derived from denoised images, and ground-truth labels using the lightweight object detection module, YOLOv8 [39]. This variant, shown in the last row of table 4, shows improved counting accuracy without sacrificing overall generation quality, suggesting that it effectively reduces the tendency to replicate objects due to textual overemphasis.

### v5

@@ Abstract
While recent advancements in generative modeling have significantly improved text-image alignment, some residual misalignment between text and image representations still remains. Some approaches address this issue by fine-tuning models in terms of preference optimization, etc., which require tailored datasets. Orthogonal to these methods, we revisit the challenge from the perspective of representation alignment—an approach that has gained popularity with the success of REPresentation Alignment (REPA) [46]. We first argue that conventional text-to-image (T2I) diffusion models, typically trained on paired image and text data (i.e., positive pairs) by minimizing score matching or flow matching losses, is suboptimal from the standpoint of representation alignment. Instead, a better alignment can be achieved through contrastive learning that leverages existing dataset as both positive and negative pairs. To enable efficient alignment with pretrained models, we propose SoftREPA—a lightweight contrastive fine-tuning strategy that leverages soft text tokens for representation alignment. This approach improves alignment with minimal computational overhead by adding fewer than 1M trainable parameters to the pretrained model. Our theoretical analysis demonstrates that our method explicitly increases the mutual information between text and image representations, leading to enhanced semantic consistency. Experimental results across text-to-image generation and text-guided image editing tasks validate the effectiveness of our approach in improving the semantic consistency of T2I generative models. Project Page: https://softrepa.github.io/.
@@ 3.1 Contrastive T2I Alignment Loss
Here, $\tau(t)$ represents time scheduling parameter. One might consider using the denoising score matching loss, $\mathbb{E}_{t,{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}]$, as a similarity measure instead of relying on its logit value. However, our experiments revealed that this formulation can introduce instability during training due to the unbounded nature of the similarity measure. To address this, SoftREPA introduce an exponential function to constrain the logit values, effectively stabilizing the training process.
@@ 3.2 The Soft Tokens
To distill the contrastive score matching objective above, we introduce a learnable soft token $\bm{s}$}, which is the only trainable parameter while pretrained model remains frozen. In particular, as illustrated in fig.˜2, we introduce a set of learnable soft tokens that vary between layers and are indexed by time $t$. The soft token at layer $k$ is defined using an embedding function with the number of soft tokens as $m$:
@@
Then, the learnable token $\bm{s}$} is optimized by minimizing $\mathcal{L}_{\text{SoftREPA}}({\bm{s}})$. During inference, the trained soft tokens are leveraged with fixed text representation from the text encoder across layers and timesteps. The detailed procedure during the image generation process based on MM-DiT architecture is described in algorithm˜1 from appendix˜B.
@@ 3.3 Relationship with Mutual Information
where the second equality follows from the definition of conditional expectation. Song et al. [36] and Kong et al. [18] further show that assuming the diffusion model is an optimal denoiser, the conditional likelihood $p_{\theta}({\bm{x}}|{\bm{y}})$ can be described as (see appendix˜A):
$p_{\theta}({\bm{x}}|{\bm{y}})=\exp(\hat{l}({\bm{x}},{\bm{y}}))\quad\text{where}\quad\hat{l}({\bm{x}},{\bm{y}})=-\frac{1}{2}\int_{0}^{T}\lambda(t)\mathbb{E}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}]dt+C.$ (16)
@@
Human Preference Text Alignment Image Quality
PM
Latency
(sec/image)
@@
Table 1: Quantitative evaluation of T2I generation with learnable soft tokens on SD1.5, SDXL, and SD3. Generation quality is evaluated on the COCO-val 5K [23] and GenEval [9] benchmark. Peak GPU memory(PM) usage and per-image latency are measured to assess computational efficiency. ImageReward, CLIP, HPS, and LPIPS are scaled by $\times 10^{2}$.
@@
Table 2: Quantitative evaluation of image editing performance with the use of soft tokens on PnP [38], MasaCtrl [4], Direct inversion [14], and FlowEdit [19]. FlowEdit is based on SD3, and others are based on SD1.5 architecture. The editing methods are evaluated on PIEBench [14], DIV2K [1], and Cat2Dog(C2D). ImageReward, CLIP, HPS, LPIPS, and SSIM are scaled by $\times 10^{2}$ and Distance is scaled by $\times 10^{3}$.
@@ 4 Experiments
Implementation details. In our experiments, to confirm the flexibility of SoftREPA, we utilize various open-source T2I diffusion models, including Stable Diffusion 1.5, Stable Diffusion XL, and Stable Diffusion 3 to evaluate text and image alignment in both image generation and text-based image editing tasks. For the training of softs token, the batch size was set to 16 with less than 30,000 iterations using two A100 GPUs. For Stable Diffusion 3, we set the length of the soft token to 4 and used 5 layers to attach the soft tokens for further experiments. In the case of Stable Diffusion 1.5 and Stable Diffusion XL, soft tokens are applied only to the Down or Middle block layers of the UNet. Refer to further implementation details in appendix˜B.
@@
We conducted text-to-image generation experiments on SD1.5, SDXL, and SD3, training soft tokens using the text-image paired COCO dataset [23]. The implementation details are provided in appendix˜B. To evaluate the generated images, we assessed human preference scores [45, 17], text-image alignment [48, 44], and image quality [47, 34] on COCO-val 5K dataset. Furthermore, we conducted evaluation on GenEval [9], compared with RankDPO [16], which is fine-tuned via preference optimization. As shown in table˜1, in COCO dataset, the proposed method generally outperforms baseline approaches in both diffusion and rectified flow models regardless of the model architecture. Additionally, in GenEval, SD3 with SoftREPA mostly outperforms RankDPO [16]. The drop in the counting metric can be primarily attributed to the tendency of diffusion models to interpret text alignment in a way that encourages the generation of multiple instances of the referenced object. Further analysis and potential mitigation strategies are discussed in appendix˜D. In fig.˜3, detailed textual descriptions are more accurately reflected in the generated images, as demonstrated in the qualitative results.
@@
In table˜1, we report the peak GPU memory usage (PM) and per-image latency for models with and without soft tokens, including SD1.5, SDXL, and SD3. All models were evaluated in float16 precision. The use of soft tokens introduces no change in peak memory and results in only a negligible increase in latency.
@@
To further demonstrate the improved text-image alignment achieved by our method, we conducted text-guided image editing experiments using PnP [38], MasaCtrl [4] with DDIM [35], Direct Inversion [14], and FlowEdit [19]. FlowEdit employs SD3, while the other methods use SD1.5. For evaluation, we used PIEBench [14], DIV2K [1], and Cat2Dog datasets. For DIV2K editing dataset, we selected 800 high-quality images and generated detailed source and target captions using LLaVA [25]; dataset construction details are provided in appendix˜C. In table˜2, the quantitative results of human preference, text alignment, and structure distance show consistent enhancements with the use of soft tokens without compromising structure and background fidelity.
Considering that FlowEdit performs text-based image editing, where performance heavily depends on text-image alignment, this result indicates improved efficiency in modeling the joint text-image distribution using the proposed method. Furthermore, our method with soft tokens achieves more efficient editing by reducing the number of editing steps to 30, compared to 33 in FlowEdit. As shown in fig.˜6, incorporating soft tokens across various CFG scales consistently enhances text-image alignment and image quality. Figure˜4 further demonstrates that performance has been improved on both single and multi-concept editing compared to vanilla SD3. Additional results and details are provided in appendix˜F.
@@
We conducted an ablation study on SD3 using the COCO val dataset [23] to evaluate the effects of the number of layers using soft text tokens and token length on generation quality (fig.˜6). A value of 0 corresponds to vanilla SD3, and the number of layers is counted from the first. Comparisons were made using either a fixed token length of 8 or a fixed depth of 5 layers. Our findings indicate that incorporating soft tokens beyond layer $7$ severely degrades image generation quality. The optimal performance, aligned with human preferences (ImageReward [45]), was observed when soft tokens were applied within layers $2\sim 5$. This aligns with the architecture, where early layers handle text-image alignment and later ones enhance fidelity [27]. Notably, CLIP scores continued to improve as more layers used soft tokens, indicating stronger adherence to text prompts, while overall image quality remained stable. Supporting qualitative results are provided in appendix˜G. For token length, using more than 8 tokens negatively impacted quality, suggesting overfitting to the fine-tuning data. Shorter lengths (1–4 tokens) preserved perceptual similarity (LPIPS), while moderate lengths (4–8 tokens) significantly improved text-image alignment (CLIP and ImageReward). Qualitative examples are shown in appendix˜G.
Comparison with LoRA
To further assess the effectiveness of our design, we examined whether contrastive learning alone could achieve comparable improvements. Specifically, we conducted an ablation study replacing the soft tokens with LoRA. As shown in table˜3, the combination of soft tokens and contrastive loss consistently outperforms all other variants across evaluation metrics. This result demonstrates the strong synergy between soft tokens and contrastive learning in enhancing text–image alignment. These findings validate the core motivation of SoftREPA—to explore the text feature space for improved alignment between textual and visual representations. Additional comparisons with different LoRA ranks and layer configurations are presented in appendix˜H.
Human Preference Text Alignment Image Quality
[…]
SD3 90.89 22.50 26.26 27.97 56.87 42.47
SD3 + LoRA + Contrastive loss 97.72 22.57 26.42 28.44 70.33 42.71
SD3 + Soft tokens + Contrastive loss (Ours) 106.32 22.49 26.92 28.70 60.76 42.99
Table 3: Quantitative comparison of T2I generation (COCO-val 1K [23]) between soft tokens and LoRA on SD3. ImageReward, CLIP, HPS, and LPIPS are scaled by $\times 10^{2}$.
Complementarity with Diffusion RL methods
We examined the complementarity between SoftREPA and diffusion reinforcement learning (RL) methods. Since these approaches are not mutually exclusive, we tested whether SoftREPA can enhance diffusion RL frameworks. Specifically, we integrated pretrained soft tokens from SoftREPA into two representative methods—Diffusion-DPO [43], which leverages preference optimization, and DDPO [3], which applies policy optimization with an external reward—using a plug-and-play approach without additional training. As shown in table˜4, incorporating SoftREPA’s soft tokens consistently improves performance. This suggests strong complementarity while diffusion RL methods align outputs via preference or reward signals, SoftREPA refines internal text–image representations through contrastive learning. Combining the two effectively unites explicit preference alignment with robust representation learning to enhance generation quality.
Human Preference Text Alignment Image Quality
[…]
Diffusion-DPO 22.26 21.68 26.51 25.60 54.19 43.96
Diffusion-DPO + Ours 33.10 21.64 27.41 25.55 53.72 43.64
DDPO -8.83 21.20 26.21 23.45 54.83 43.85
DDPO + Ours 8.07 21.25 27.13 23.75 54.18 43.59
Table 4: Complementarity of SoftREPA with diffusion RL methods (Diffusion-DPO [43], DDPO [3]). Soft tokens are integrated in a plug-and-play manner during inference. Generation quality is evaluated on the COCO-val 1K dataset [23]. ImageReward, CLIP, HPS, and LPIPS are scaled by $\times 10^{2}$.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT) (No. RS-2024-00457882, AI Research Hub Project). This work was supported by the National Research Foundation of Korea under Grant RS-2024-00336454.
@@ Appendix A Information Theoretical Analysis of Diffusion Models
$\log p({\bm{x}})=\frac{1}{2}\int_{0}^{T}\lambda(t)\mathbb{E}_{{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\|^{2}]dt+C$ (21)
@@
$\log p({\bm{x}}|{\bm{y}})=\frac{1}{2}\int_{0}^{T}\lambda(t)\mathbb{E}_{{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}]dt+C$ (22)
@@ Appendix B Implementation Details
The hyperparameters used during training are listed in table˜5. For Stable Diffusion 3, the soft tokens are optimized solely using the contrastive score matching loss ($L_{\text{SoftREPA}}$). In contrast, for Stable Diffusion 1.5 and Stable Diffusion XL, optimization combines the contrastive score matching loss ($L_{\text{SoftREPA}}$) with a small weighting of the denoising score matching loss ($L_{\text{DSM}}$).table˜6 shows that the model consistently achieves improved performance regardless of whether the denoising score matching loss is applied. We observed that initializing the soft tokens with a random distribution led to performance degradation in Stable Diffusion 1.5. To address this, we initialized the soft tokens using unconditional text embeddings. The code is publicly available at https://github.com/softrepa/SoftREPA.
@@
Human Preference Text Alignment Image Quality
[…]
SD1.5 - 17.72 21.47 26.45 25.08 24.59 43.80
$L_{\text{SoftREPA}}+L_{\text{DSM}}$ 32.89 21.50 27.33 25.18 23.43 43.38
$L_{\text{SoftREPA}}$ 32.63 21.61 27.10 25.94 29.25 44.09
SDXL - 75.06 22.38 26.76 27.35 24.69 42.05
$L_{\text{SoftREPA}}+L_{\text{DSM}}$ 85.29 22.62 26.80 28.30 26.04 42.39
$L_{\text{SoftREPA}}$ 81.09 22.59 26.87 28.32 26.42 42.69
Table 6: Quantitative comparison of T2I generation between contrastive score matching loss ($L_{\text{SoftREPA}}$) with and without denoising score matching loss ($L_{\text{DSM}}$). Generation quality is evaluated on the COCO-val 5K [23]. ImageReward, CLIP, HPS, and LPIPS are scaled by $\times 10^{2}$.
@@
A detailed description of the image generation algorithm on MM-Dit is provided in algorithm˜1. At each layer, a distinct set of soft tokens is prepended to the text features and used exclusively within that layer. These soft tokens are not carried over to subsequent layers; instead, new soft tokens are introduced or omitted as appropriate. Their primary role is to guide the text tokens toward better text-image alignment, particularly during the early stages of the model.
@@
For the memory efficiency metrics reported in table˜1, values were measured by averaging results over 50 runs with an A100 GPU. Image resolution was set to $512\times 512$ for SD1.5 and $1024\times 1024$ for SDXL and SD3.
Algorithm 1 Image Generation with Soft Tokens in $M$-DiT
@@
4:Number of layers NN, Threshold layer $L$
@@
18:    $\bm{H}_{\text{text}}^{(k,t)}\leftarrow\hat{\bm{H}}_{\text{text}}^{(k,t)}[-n:,:]$ $\triangleright$ Drop soft tokens
@@ Appendix D Discussion on counting metric
As shown in section˜4, our model demonstrates notable improvements in generating single and multiple objects, as well as in color and position-aware synthesis. However, it exhibits a relative decline in accurately generating the specified number of objects described in the text prompts. We hypothesize that this drop in the counting metric stems from the soft tokens excessively emphasizing textual cues, which can lead to overgeneration of object instances. To address this, we limited the use of soft tokens to the early layers of the model. As presented in table˜7, applying soft tokens only to layers $1\sim 2$ preserves strong overall performance while mitigating the degradation in counting accuracy observed when soft tokens are applied across layers $1\sim 5$.
@@
To further enhance counting fidelity, we trained SoftREPA with a lightweight object-counting loss. Specifically, we employ a mean squared error (MSE) loss between the predicted object count, which is derived from denoised images, and ground-truth labels using the lightweight object detection module, YOLOv8 [40]. This variant, shown in the last row of table˜7, shows improved counting accuracy without sacrificing overall generation quality, suggesting that it effectively reduces the tendency to replicate objects due to textual overemphasis.